Please help me solve this problem. It is a-3/5=-7/10.
Thank you so much!

Answer by stanbon(57352)

a-3/5=-7/10
-----
Multiply thru by 10 to get:
10a -6 = -7
10a = -1
a = -1/10
=============
Cheers,
Stan H.
=============